Find the characteristic equation of the matrix

A= [ 1 2 3
4 1 -1
-3 2 5 ]
and verify Cayley-Hamilton theorem for it.
1
Expert's answer
2020-12-21T17:01:25-0500

The characteristic equation of the given matrix "A" is "det(A-\lambda I)=0", where "\lambda" is any scalar.

"\Rightarrow \left | 
 \begin{matrix}
 1-\lambda & 2 & 3 \\
 4 & 1-\lambda & -1 \\
 -3 & 2 & 5-\lambda 
 \end{matrix}
\ \right |=0"

"\Rightarrow (1-\lambda )\left [ (1-\lambda )(5-\lambda )+2 \right ]-2\left [ 4(5-\lambda )-3 \right ]+3\left [ 8+3(1-\lambda ) \right ]=0"

"\Rightarrow (1-\lambda )(\lambda ^{2}-6\lambda +7)-2(17-4\lambda )+3(11-3\lambda )=0"

"\Rightarrow -\lambda ^{3}+7\lambda ^{2}-14\lambda +6=0"

Therefore, the characteristic of the given matrix is given by

"\lambda ^{3}-7\lambda ^{2}+14\lambda -6=0"

Now let us verify the Cayley- Hamilton theorem.

The matrix equation is given by

"A ^{3}-7A ^{2}+14A -6I=O"

"A^2=\left [ 
 \begin{matrix}
 1 & 2 & 3 \\
 4 & 1 & -1 \\
 -3 & 2 & 5
 \end{matrix}
\ \right ]\left [ 
 \begin{matrix}
 1 & 2 & 3 \\
 4 & 1 & -1 \\
 -3 & 2 & 5
 \end{matrix}
\ \right ]=\left [ 
 \begin{matrix}
 0 & 10 & 16 \\
 11 & 7 & 6 \\
 -10 & 6 & 14
 \end{matrix}
\ \right ]"

And

"A^3=\left [ 
 \begin{matrix}
 0 & 10 & 16 \\
 11 & 7 & 6 \\
 -10 & 6 & 14
 \end{matrix}
\ \right ]\left [ 
 \begin{matrix}
 1 & 2 & 3 \\
 4 & 1 & -1 \\
 -3 & 2 & 5
 \end{matrix}
\ \right ]=\left [ 
 \begin{matrix}
 -8 & 42 & 70 \\
 21 & 41 & 56 \\
 -28 & 14 & 34
 \end{matrix}
\ \right ]"

Now

"A ^{3}-7A ^{2}+14A -6I"

"\left [ 
 \begin{matrix}
 -8 & 42 & 70 \\
 21 & 41 & 56 \\
 -28 & 14 & 34
 \end{matrix}
\ \right ]-\left [ 
 \begin{matrix}
 0 & 70 & 112 \\
 77 & 49 & 42 \\
 -70 & 42 & 98
 \end{matrix}
\ \right ]"

"+\left [ 
 \begin{matrix}
 14 & 28 & 42 \\
 56 & 14 & -14 \\
 -42 & 28 & 70
 \end{matrix}
\ \right ]-\left [ 
 \begin{matrix}
 6 & 0 & 0 \\
 0 & 6 & 0 \\
 0 & 0 & 6
 \end{matrix}
\ \right ]"


"=\left [ 
 \begin{matrix}
 0 & 0 & 0 \\
 0 & 0 & 0 \\
 0 & 0 & 0
 \end{matrix}
\ \right ]"

"=O"

Therefore, "A ^{3}-7A ^{2}+14A -6I=O"

Hence, Cayley-Hamilton Theorem is verified.